# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v3

@@ 3.2 Extending explainability methods in 3D space
In this study, we used two distinct explainable techniques: the widely-utilized sensitivity local explainability technique in medical imaging applications known as the GradCam method (Selvaraju et al. (2017)), and a robust attribution explainability technique called SHAP (Lundberg and Lee (2017)). Previously emphasized is the significance of the three-dimensional space in neuroscience applications. Consequently, there arises a necessity to extend these 2D XAI methods into three dimensions. In the pursuit of computing the class-discriminative localization map encompassing width $w$, height $h$, and depth $d$ within a specific 3D brain MR image corresponding to a class $c$ (PCS or noPCS), the computation involves determining the gradient of the score for class $c$, denoted as $y^{c}$, in relation to the $o^{th}$ feature activation map ($A^{o}$) of the final convolution layer in each deep network. To determine the importance weights ($\alpha^{c}_{o}$) for each $o$ feature activation map, global average pooling is employed over the width ($i$), height ($j$), and depth ($k$) of each feature.
$\alpha^{c}_{o}=\frac{1}{Z}\sum_{i}{{\sum_{j}\sum_{k}\frac{dy^{c}}{dA_{ijk}^{o}}}}$ (4)
@@
$GradCam=ReLU(\sum_{o}{\alpha^{c}_{o}A_{ijk}^{o}})$ (5)
SHAP (SHapley Additive exPlanations; Lundberg and Lee (2017)) computes the attribution of each pixel of an input image for a specific prediction of a computer vision task. The attribution explainability methods follow the definition of additive feature attribution mainly as a linear function of:
@@
where $f$ is the prediction network, $g(f,x)$ is the explanation model, $\phi_{i}$ is the importance of each feature attribution ($\phi_{i}\in\mathbb{R}$), and $M$ is the number of simplified input features (pixels). Shapley value estimation is one of the main mathematical formulations that the SHAP algorithm used to assign an importance value to each feature, representing the effect on the model prediction of including that feature (attribution). If we define a subset $S$ of the total feature space ($F$) of an input 3D image $I$ ($i=1...N$, where $N$ is the number of samples in the dataset), and $x_{i}$ is a 3D matrix of width $w$ and height $h$ and depth $d$ for the $i^{th}$ sample, and $x_{S}$ is the subset of chosen features in the 3D space, then:
@@
Here, $f_{S(i)}$ is a model trained with the presented $x_{s}$ features, and $f_{S}$ is another model trained with the features withheld. For our study, we used Deep SHAP Lundberg and Lee (2017) to describe our deep learning network models. In this approach, we used a chain rule and linear approximation as described in Lundberg and Lee (2017).
@@ 3.3 The XAI 3D-Framework
The main purpose of the XAI 3D-Framework is to provide global explanations that can lead to new pattern discovery. To achieve this, we used two XAI methods, GradCam and SHAP in 3D space to extract local explanations from the deep learning classifiers (simple-3D-CNN, simple-3D-MHL; Fig. 2 b.). Concurrently, we assigned faithfulness and complexity scores to evaluate the quality of the local explanations, ensuring their effectiveness reaches an acceptable threshold (see Supplementary material Table 2). Following that, we used Principal Component Analysis (PCA; Pearson (1901)) to derive global explanations from the two XAI methods. Moreover, PCA dimensionality reduction was applied to both the sulcal skeleton and the grey/white surface 3D brain inputs to identify a variety of global feature importances (feature-shape; see Fig. 3). We experimented with a range of different PCA number of components, to find the most optimal configuration in terms of time efficiency and the proportion of cohort variance captured (more than 80.0%). The optimal solution was found with six components. Based on a preliminary research, we found that the global explanations derived from each of the two XAI methods (GradCam and SHAP), yield better results in terms of faithfulness and complexity when we average the six components of the PCA analysis using a weighted tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001] (see Supplementary Material Table 2; PCA0-SHAP vs total-SHAP, and PCA0-GradCam vs total-GradCam). Consequently, to generate the final global explanation represented from the GradCam (total-GradCam) and SHAP (total-SHAP), we apply the weighted average to the six PCA components (see Fig. 3). The same approach is applied in the feature-shape domain to determine the global statistical feature importance of the sulcal skeleton and the grey/white surface in the 3D brain inputs (total-Shape; see Fig. 3). The weighted average formulation is defined as follows:
@@
Figure 2: Classification task determination and the transition from local to global 3D explaination. a, Illustration of the no paracingulate sulcus condition (no PCS), and the paracingulate sulcus condition (PCS, in green line) on two left hemispheres 3D white matter reconstructions obtained with BrainVISA. The cingulate sulcus is coloured in yellow and blue and the callosal sulcus is coloured in purple. b, The 3D explainable framework that provides both local and global interpretations and explanations of our deep learning 3D classification network’s results. The ratio of the faithfulness and complexity metrics are compute in that stage. In this example we include only the GradCam explainability method just for simplicity.
@@ 3.4 The global explanation of 3D-Framework
To compute the global explanation for the 3D framework, we first manually align and rescale the two global XAI explanations (total-SHAP, total-GradCam) to the global statistical features (total-Shape). Then, we use equation 8, incorporating a three-component weight tensor of $[0.85,0.5,0.1]$. We conducted an ablation study to identify the optimal combination of the weight tensor. The different cases we examined were denoted as 851, 815, 185, 158, 518, and 581, which represent a fixed-order representations of the methods in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations. The weights correspond to 0.85 as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For example, 3D-Framework-851 refers to the proposed 3D-Framework with weight values of 0.85 for total-Shape, 0.5 for total-SHAP, and 0.1 for total-GradCam. These components were derived from both the sulcal skeleton and the grey/white surface inputs of the total TOP-OSLO cohort, considering both the right and left hemispheres.
To minimize potential bias that might occur by relying on only one deep learning network, we applied this approach to both the simple-3D-CNN and simple-3D-MHL networks. This allowed us to identify the most relevant features for each classification class by examining the overlap between the results of the two models. Identifying these features provides insights into the mechanisms driving the network’s decision-making process. For enhanced clarity, the key brain sub-regions of interest, corresponding to the sulcal skeleton and grey/white surface, are visually depicted in Fig. 1a. Finally, the global explanation from the 3D framework was registered on a sulcal probabilistic atlas to illustrate the model’s pattern for determining the presence or absence of the paracingulate sulcus. The registration process involved affine transformations, including translation, rotation, scaling, and geometry adjustments, to align with the probabilistic atlas.
@@ 4.1 Cohort’s description and pre-processing image analysis
We used the structural MRI of 596 participants from the TOP-OSLO study Mørch-Johnsen et al. (2017) for a binary classification task. The participants encompassed individuals within the unaffected control spectrum (262), those within the schizophrenia spectrum (183), and those within the bipolar disorder spectrum (151). The participants were imaged using a 1.5 T Siemens Magnetom Sonata scanner (Siemens Medical Solutions, Erlangen, Germany). The modality used was T1-weighted structural MRI, and the input files consisted of DICOM images. Two experts, A.A. and H.V., performed image annotations, categorizing them into two classes: ’no paracingulate sulcus’ (noPCS) and ’paracingulate sulcus’ (PCS). This annotation was based on the protocol described in Mitchell et al. (2023), with further details available in the supplementary materials under ’1.1 PCS classification of TOP-OSLO’ and it is illustrated in Fig. 2 a. Image analysis techniques were applied to all slices to reduce artefacts. We used noise filters such as binomial deconvolution, Landweber deconvolution Vonesch and Unser (2008), and curvature anisotropic diffusion image filters Perona and Malik (1990) to reduce noise in the images.
Figure 3: The proposed global 3D-Framework explanation. A weighted averaging (Weight tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001]) of six PCA components produces the Global PCA image for PCA-Shape, PCA-GradCam, and PCA-SHAP. Following this, a weighted averaging (Weight tensor: [0.85, 0.5, 0.1]) of the three Global PCA overlapping images extracts the total overlapping image. This total overlapping image is then registered on a sulcal probabilistic atlas to unveil the model’s pattern for determining the presence or absence of the paracingulate sulcus.
@@ 4.2 Processing of two distinct images inputs
We initially processed the brain structural MRIs using the BrainVISA software Cointepas et al. (2001) and extracted two different images for our input: the grey-white surface and the sulcal skeleton. These were extracted from the raw MRI from a succession of bias correction, histogram analysis, brain segmentation, hemisphere separation, dichotomization of the white matter from the union of grey matter and cerebrospinal fluid, and skeletonization of the result, as detailed in Rivière et al. (2002). Specifically, the grey/white boundary was obtained by minimizing a Markov field and the segmentation used homotopic deformations of the hemisphere bounding box, resulting in the grey-white surface, where voxels are dichotomised into either grey or white. The skeleton was then derived from this object by applying a homotopic erosion embedding a watershed alogrithm, that preserves the initial topology, resulting in the sulcal skeleton. These two modalities were then used to train and evaluate our networks as well as our explainability methods. Fig. 1a,b,c. show the structural MRI and the corresponding grey-white surface and sulcal skeleton outputs from BrainVISA.
4.3 Hyper-parameters initialization
After randomly shuffling the data, each dataset was split into training, validation, and testing sets containing $70\%$, $20\%$, and $10\%$ of the total number of images, respectively. Sparse categorical cross-entropy was used as the cost function and the loss function was optimized using the Adam algorithm Kingma and Ba (2014). The learning rate was exponentially decreased during the first 50 epochs and then fixed at $0.0001$ for the last 50 epochs. To train the networks, an early stopping criterion of 10 consecutive epochs was employed and a maximum of 100 epochs was used for all input modalities (sulcal skeleton and grey/white matter) for both the left (L) and right (R) hemispheres. Finally, we use data augmentation techniques including rotation (around the center of the image by a random angle from the range $[-15^{\circ},15^{\circ}]$), width shift (up to $20$ pixels), height shift (up to $20$ pixels), and Zero phase Component Analysis (ZCA; Bell and Sejnowski (1997))whitening (add noise in each image) to avoid overfitting.
@@ 4.3 Hyper-parameters initialization
A crucial aspect of this study lies in evaluating how accurate and comprehensive is a local or global explanation. To derive a useful explanation, two primary scores play a pivotal role: faithfulness and complexity. An intuitive way to assess the quality of an explanation is by measuring its ability to accurately capture how the predictive model reacts to random perturbations Yeh et al. (2019). For a deep neural network $f$, and input features $\boldsymbol{x}$, the feature importance scores (also known as "attribution scores") are derived in a way such that when we set particular input features $\boldsymbol{x_{s}}$ to a baseline value $\boldsymbol{x_{s}^{f}}$, the change in the network’s output should be proportional to the sum of attribution scores of the perturbed features $\boldsymbol{x_{s}}$. We quantify this by using the Pearson correlation between the sum of the attributions of $\boldsymbol{x_{s}}$ and the difference in the output when setting those features to a reference baseline Bhatt et al. (2020). Thus, we define the faithfulness of an explanation method $g$ as:
@@
If for an image $\boldsymbol{x}$, explanation $g$ highlights all $d$ features, then it may be less comprehensible and more complex than needed (especially for large $d$). It is important to compute the level of complexity, as an efficient explanation has to be maximally comprehensible Bhatt et al. (2020). If $P_{g}$ is a valid probability distribution and the $P_{g}(i)$ is the fractional contribution of feature $x_{i}$ to the total magnitude of the attribution, then we define the complexity of the explanation $g$ for the network $f$ as:
@@
In order to evaluate these two explainability metrics, we used the software developed by Hedström et al. (2023). This software package is a comprehensive toolkit that collects, organizes, and evaluates a wide range of performance metrics, proposed for explanation methods. We note that we used a zero baseline (’black’; $\boldsymbol{x_{s}^{f}}$= $\boldsymbol{0}$), and 70 random perturbations to calculate the Faithfulness score.
Finally, to extract the faithfulness and complexity scores of the global explanations for the total-SHAP, total-GradCam, and the proposed 3D-Framework, we utilized the software developed by Hedström et al. (2023). As a first step we manually align and rescale the two global XAI explanations (total-SHAP and total-GradCam) to the global statistical features (total-Shape). The input image consisted of the total-Shape results, while the total-GradCam, total-SHAP, and 3D-Framework global explanations served as reference explanations, respectively. In this context, the score of a global explanation makes sense, as the individual inputs brains of the cohorts are aligned in the same template, and the most significant variability of each class is assigned in the total-Shape. Consequently, we anticipate the classifier to predict correctly whether a patient has a PCS or not, as observed.
@@ 4.4 Data availability
This study used the datasets of the TOP-OSLO Mørch-Johnsen et al. (2017) which can be obtained from University of Oslo upon request, subject to a data transfer agreement.

### v4

@@ 3.2 Extending explainability methods in 3D space
We used two distinct explainable techniques: the widely-utilized sensitivity local explainability technique in medical imaging applications known as the GradCam method ((Selvaraju et al. (2017)), and a robust attribution explainability technique called SHAP ((Lundberg and Lee (2017)).
The significance of the 3D space in neuroscience applications emphasises a necessity to extend 2D XAI methods into three dimensions. In the pursuit of computing the class-discriminative localization map encompassing width $w$, height $h$, and depth $d$ within a specific 3D brain MRI corresponding to a class $c$ (PCS or noPCS), the computation involves determining the gradient of the score for class $c$, denoted as $y^{c}$, in relation to the $n^{th}$ feature activation map ($A^{n}$) of the final convolution layer in each deep network. To determine the importance weights ($\alpha^{c}_{n}$) for each $n$ feature activation map, global average pooling is employed over the width ($i$), height ($j$), and depth ($k$) of each feature.
$\alpha^{c}_{n}=\frac{1}{Z}\sum_{i}{{\sum_{j}\sum_{k}\frac{dy^{c}}{dA_{ijk}^{n}}}}$ (4)
@@
$GradCam=ReLU(\sum_{n}{\alpha^{c}_{n}A_{ijk}^{n}})$ (5)
SHAP computes the attribution of each pixel of an input image for a specific prediction of a computer vision task. Attribution explainability methods follow the definition of additive feature attribution mainly as a linear function of:
@@
where $f$ is the prediction network, $g(f,x)$ is the explanation model, $\phi_{i}$ is the importance of each feature attribution ($\phi_{i}\in\mathbb{R}$), and $M$ is the number of simplified input features (pixels). Shapley value estimation is one of the main mathematical formulations that the SHAP algorithm uses to assign an importance value to each feature, representing the effect on the model prediction of including that feature (attribution). If we define a subset $S$ of the total feature space ($F$) of an input 3D image ($i=1...N$, where $N$ is the number of samples in the dataset), and $x_{i}$ is a 3D matrix of width $w$ and height $h$ and depth $d$ for the $i^{th}$ sample, and $x_{S}$ is the subset of chosen features in the 3D space, then:
@@
Here, $f_{S(i)}$ is a model trained with the presented $x_{s}$ features, and $f_{S}$ is another model trained with the features withheld. For our study, we used Deep SHAP (Lundberg and Lee (2017) to describe our deep learning network models. This approach uses a chain rule and linear approximation as described in (Lundberg and Lee (2017)).
@@ 3.3 The XAI 3D-Framework
The main purpose of the XAI 3D-Framework is to provide global explanations that can lead to the discovery of new patterns. To achieve this, we used two XAI methods, GradCam and SHAP in 3D space to extract local explanations from the deep learning classifiers (simple-3D-CNN, simple-3D-MHL; Fig. 2 b.). Concurrently, we assigned faithfulness and complexity scores to evaluate the quality of the local explanations ensuring that their effectiveness reached an acceptable threshold (see Supplementary material Table 2). Following that, we used Principal Component Analysis (PCA; (Pearson (1901)) to derive global explanations from the two XAI methods. PCA dimensionality reduction was applied to both the sulcal skeleton and the grey-white surface 3D brain inputs (analytical description of the 3D inputs please advice sub-section 4.2) to identify a variety of global feature importance (feature-shape; see Fig. 3). We experimented with a range of different numbers of PCA components to find the optimal configuration in terms of time efficiency and the proportion of cohort variance captured ($>80.0\%$). The optimal solution was found with six components.
Based on preliminary research, we found that the global explanations derived from each of the two XAI methods (GradCam and SHAP) yielded better results in terms of faithfulness and complexity when we averaged the six components of the PCA analysis using a weighted tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001] (see Supplementary Material Table 2; PCA0-SHAP vs total-SHAP, and PCA0-GradCam vs total-GradCam). Consequently, to generate the final global explanation represented from the GradCam (total-GradCam) and SHAP (total-SHAP), we applied the weighted average to the six PCA components (see Fig. 3). The same approach was applied in the feature-shape domain to determine the global statistical feature importance of the sulcal skeleton and the grey-white surfaces in the 3D brain inputs (total-Shape; see Fig. 3). The weighted average formulation is defined as follows:
@@
Figure 2: Classification task determination and the transition from local to global 3D explaination. a, Illustration of the no PCS condition (no PCS), and the PCS condition (PCS, in green line) on two left hemisphere 3D white matter reconstructions obtained with BrainVISA. The cingulate sulcus is coloured in yellow and blue and the callosal sulcus is coloured in purple. b, The 3D explainable framework that provides both local and global interpretations and explanations of our deep learning 3D classification network’s results. The ratio of the faithfulness and complexity metrics were computed at that stage. In this example we include only the GradCam explainability method for simplicity.
@@ 3.4 The global explanation of 3D-Framework
To compute the global explanation for the 3D framework, we first manually aligned and rescaled the two global XAI explanations (total-SHAP, total-GradCam) to the global statistical features (total-Shape). Then, we used equation 8, incorporating a three-component weight tensor of $[0.85,0.5,0.1]$. We conducted an ablation study to identify the optimal combination of the weight tensor. The different cases we examined were denoted as 851, 815, 185, 158, 518, and 581, which are fixed-order representations of the methods in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations. The weights correspond to 0.85 as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For example, 3D-Framework-851 refers to the proposed 3D-Framework with weight values of 0.85 for total-Shape, 0.5 for total-SHAP, and 0.1 for total-GradCam. These components were derived from both the sulcal skeleton and the grey-white surface inputs of the total dataset, considering both the right and left hemispheres.
To minimize potential bias that might occur by relying on only one deep learning network, we applied this approach to both the simple-3D-CNN and simple-3D-MHL networks. Identifying the significant features of the networks explanations provides insights into the mechanisms driving the network’s decision-making process. For enhanced clarity, the key brain sub-regions of interest, corresponding to the sulcal skeleton and grey-white surface, are visually depicted in Fig. 1a. Finally, the global explanation from the 3D framework was registered on a sulcal probabilistic atlas to illustrate the model’s pattern associated with determining the presence or absence of the PCS. The registration process involved affine transformations, including translation, rotation, scaling, and geometry adjustments, to align with the probabilistic atlas.
@@ 4.1 Cohort’s description and pre-processing image analysis
We used the structural MRI of 596 participants from the TOP-OSLO study (Mørch-Johnsen et al. (2017)) for a binary classification task. The participants encompassed individuals with a diagnosis on the schizophrenia spectrum (183), on the bipolar disorder spectrum (151), and unaffected control participants (262). T1-weighted images were acquired using a 1.5 T Siemens Magnetom Sonata scanner (Siemens Medical Solutions, Erlangen, Germany).
Two experts, A.A. and H.V., performed image annotations, categorizing them into two classes: ’no paracingulate sulcus’ (noPCS) and ’paracingulate sulcus’ (PCS). This annotation was based on the protocol described in (Mitchell et al. (2023)) with further details available in the supplementary materials under ’1.1 PCS classification of TOP-OSLO’ and is illustrated in Fig. 2a.
@@ 4.2 Processing of two distinct images inputs
We initially processed the brain structural MRIs using the BrainVISA software (Cointepas et al. (2001)) extracting two images as inputs for the classifier: the grey-white surface and the sulcal skeleton. These were extracted from the raw MRI with an established protocol consisting of bias correction, histogram analysis, brain segmentation, hemisphere separation, dichotomization of the white matter from the union of grey matter and cerebrospinal fluid, and skeletonization of the result, as detailed in (Rivière et al. (2002)). Specifically, the grey-white surface was obtained by minimizing a Markov field and the segmentation used homotopic deformations of the hemisphere bounding box, resulting in the grey-white surface, where voxels are dichotomised into either grey or white. The skeleton was then derived from this object by applying a homotopic erosion embedding a watershed alogrithm that preserves the initial topology resulting in the sulcal skeleton. These two modalities were then used to train and evaluate our networks as well as for our explainability methods. Fig. 1a,b,c. show the structural MRI and the corresponding grey-white surface and sulcal skeleton outputs from BrainVISA.
Figure 3: The proposed global 3D-Framework explanation. A weighted averaging (Weight tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001]) of six PCA components produces the Global PCA image for PCA-Shape, PCA-GradCam, and PCA-SHAP. Following this, a weighted averaging (Weight tensor: [0.85, 0.5, 0.1]) of the three Global PCA overlapping images extracted the total overlapping image. This total overlapping image was then registered on a sulcal probabilistic atlas to unveil the model’s pattern for determining the presence or absence of the PCS.
4.3 Hyper-parameter initialization
After randomly shuffling the data, each dataset was split into training, validation, and testing sets containing $70\%$, $20\%$, and $10\%$ of the total number of images, respectively. Sparse categorical cross-entropy was used as the cost function and the loss function were optimized using the Adam algorithm (Kingma and Ba (2014)). After manual hyper-parameter searching the best learning rate was a value of $0.0001$. We utilized a strategy of exponentially decreased during the first 50 epochs and then fixed at $0.0001$ for the last 50 epochs. To train the networks, an early stopping criterion of 10 consecutive epochs was employed and a maximum of 100 epochs was used for both input modalities (sulcal skeleton and grey-white surface) for both the left (L) and right (R) hemispheres. Finally, we use data augmentation techniques including rotation (around the center of the image by a random angle from the range $[-15^{\circ},15^{\circ}]$), width shift (up to $20$ pixels), height shift (up to $20$ pixels), and Zero phase Component Analysis (ZCA; Bell and Sejnowski (1997)) whitening (add noise in each image) to avoid overfitting.
@@ 4.3 Hyper-parameter initialization
A crucial aspect of this study lies in evaluating how accurate and comprehensive were the local and global explanations. To derive a useful explanation, two primary scores play a pivotal role: faithfulness and complexity. An intuitive way to assess the quality of an explanation is by measuring its ability to accurately capture how the predictive model reacts to random perturbations (Yeh et al. (2019)). For a deep neural network $f$, and input features $\boldsymbol{x}$, the feature importance scores (also known as "attribution scores") were derived in a way such that when we set particular input features $\boldsymbol{x_{s}}$ to a baseline value $\boldsymbol{x_{s}^{f}}$, the change in the network’s output was proportional to the sum of attribution scores of the perturbed features $\boldsymbol{x_{s}}$. We quantified this by using the Pearson correlation between the sum of the attributions of $\boldsymbol{x_{s}}$ and the difference in the output when setting those features to a reference baseline (Bhatt et al. (2020)). Thus, we define the faithfulness of an explanation method $g$ as:
@@
If for an image $\boldsymbol{x}$, explanation $g$ highlights all $d$ features, then it may be less comprehensible and more complex than needed (especially if $d$ is large). It is important to compute the level of complexity, as an efficient explanation has to be maximally comprehensible (Bhatt et al. (2020)). If $P_{g}$ is a valid probability distribution and the $P_{g}(i)$ is the fractional contribution of feature $x_{i}$ to the total magnitude of the attribution, then we define the complexity of the explanation $g$ for the network $f$ as:
@@
In order to evaluate these two explainability metrics, we used the software developed by (Hedström et al. (2023). This software package is a comprehensive toolkit that collects, organizes, and evaluates a wide range of performance metrics, proposed for explanation methods. Note that we used a zero baseline (’black’; $\boldsymbol{x_{s}^{f}}$= $\boldsymbol{0}$) and 70 random perturbations to calculate the faithfulness score.
Finally, to extract the faithfulness and complexity scores of the global explanations for the total-SHAP, total-GradCam, and the proposed 3D-Framework, we utilized again the software developed by (Hedström et al. (2023)). As a first step we manually aligned and rescaled the two global XAI explanations (total-SHAP and total-GradCam) to the global statistical features (total-Shape). The input image consisted of the total-Shape results, while the total-GradCam, total-SHAP, and 3D-Framework global explanations served as reference explanations, respectively. In this context, the score of a global explanation makes sense, as the individual input brains of the cohort were aligned in the same template and the most significant variability of each class was assigned in the total-Shape. Consequently, we anticipated the classifier to classify correctly whether a MRI has a PCS or not.
@@ 4.4 Data availability
This study used the dataset of the TOP-OSLO cohort (Mørch-Johnsen et al. (2017)) which can be obtained from University of Oslo upon request, subject to a data transfer agreement.